OpenStudy (felecia):

Evaluate 8x + 7 given that x - 3 = 10.

OpenStudy (kittiwitti1):

x-3=10 get x alone on one side then put the x value you get into (8x+7) and solve

OpenStudy (falcon):

x-3 = 10 so x = 10+3 so x = 13 Put this value in 8x + 7 8*13 + 7 =111

OpenStudy (kittiwitti1):

\[\huge{↑}\]\(\Large\text{The previous response above is a direct answer.}\) Direct answers are against the rules of the site.